# Optimal Kron-based Reduction of Networks (Opti-KRON) for Three-phase Distribution Feeders

### v1

@@ I Introduction
The rapid integration of distributed energy resources (DERs) and advanced control schemes has significantly increased the complexity of distribution networks, posing challenges for optimization and real-time decision-making [1]. Detailed network models with thousands of nodes are computationally intensive and are impractical for large-scale optimal power flow and control problems. Consequently, research on network reduction has emerged, where proposed network reduction tools seek to simplify network models while preserving essential electrical characteristics.
@@
One network reduction technique is bus elimination via Kron reduction [2], where selected nodes are eliminated via the Schur complement of the nodal admittance matrix, resulting in an equivalent network that preserves the impedance seen between retained nodes. However, the set of reducible nodes is restricted to those with zero current injection.  [3] and [4] develop methods for network reduction on distribution systems, but both are limited to single-phase networks and do not extend to real-world unbalanced three-phase feeders. References [5, 6, 7], and [8] propose methods for network reduction on distribution systems considering multiphase unbalanced models. Nevertheless, a common limitation across these studies lies in their dependence on a predefined set of nodes that must be kept. Whether these nodes are specified as critical nodes or boundary nodes, this requirement restricts flexibility and may reduce the accuracy of the reduced networks.
@@
In prior works [9] and [10], an optimal Kron-based reduction of networks (Opti-KRON) framework has been developed for single-phase balanced feeders, where the reduced networks do not rely on any predefined set of nodes and explicitly trades off reduction level against accuracy via tunable bounds on voltage magnitude error. A mixed-integer linear program (MILP) was proposed to identify the optimal set of nodes and clusters that maximize reduction while minimizing voltage deviations across representative loading scenarios. In addition, because Kron reduction can yield dense, meshed networks [2][9]), the framework introduced a radialization step to ensure the reduced network remains radial. However, the previous Opti-KRON formulation is limited to single-phase networks. As a result, no existing optimal network reduction method can efficiently reduce unbalanced three-phase distribution systems.
@@
In this paper, we extend the Opti-KRON framework of [9] to unbalanced three-phase distribution networks. We develop a three-phase Kron-based reduction and radialization procedure. To formulate the optimal reduction, we first present that generalizes the single-phase Opti-KRON to three-phase feeders. Furthermore, we introduce an exhaustive-search algorithm that serves as a benchmark to the MILP formulation. The proposed exhaustive search offers several advantages. It operates directly on complex values without requiring decomposition to rectangular coordinates. Second, it can evaluate nonconvex objective functions. Third, its structure supports parallel computation on GPUs. Consequently, the proposed method achieves significant computational speedups on large-scale feeders while increasing reduction accuracy. This paper makes the following contributions:
@@
The remainder of this paper is organized as follows. In Section II, the three-phase network model and Kron reduction are summarized. Section IV introduces our exhaustive-search framework for optimal Kron-based reduction. We then describe the radialization step in Section V. Experimental results on large unbalanced feeders appear in Section VI. Section VII concludes with a summary and future directions.
@@ II-A Network Representation
Consider a distribution network as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n=\lvert\mathcal{V}\rvert$ nodes and $m=\lvert\mathcal{E}\rvert$ branches. We define the network topology via the adjacency matrix $\Lambda\in\{0,1\}^{n\times n}$, where $\Lambda_{ij}=1$, if a branch connects nodes $i$ and $j$, and $\Lambda_{ij}=0$ otherwise. Each node $i$ carries a subset of phases $\phi(i)\subseteq\{a,b,c\}$. Each node $i$ is associated with a three-phase voltage vector $V_{i}\in\mathbb{C}^{3}$ and current injection vector $I_{i}\in\mathbb{C}^{3}$. Complex nodal voltage and current injection vectors are defined as
@@ II-B Kron Reduction of Three-Phase Networks
The Kron reduction of $Y$ is the Schur complement of $Y_{rr}$ [2]. Thus, the Kron reduced admittance matrix, denoted as $Y_{\rm Kron}$, is given by
@@
where $\{\cdot\}^{+}$ represents Moore-Penrose Pseudoinverse  [^2] The pseudo-inverse is utilized to account for zero rows and columns associated with missing phases during elimination. and
@@ II-C Clustering for Kron-based Reduction
As shown in (8) and (10), Kron reduction preserves voltage profile when the reduced nodes have zero current injection. To enable reduction beyond strictly zero-injection nodes, we introduce an assignment scheme, where each reduced node moves its current injection to a kept adjacent node. We refer to the kept nodes $\mathcal{K}$ as super-nodes and they contain the aggregated current of themselves and the reduced nodes assigned to them. Each super-node $k\in\mathcal{K}$ represents a cluster consists of itself and its assigned children-nodes. These assignments, however, must obey several rules:
@@
    Phase availability: a reduced node $i$ can only be assigned to a super-node $k$ if $\phi(i)\subseteq\phi(k)$.
@@
Now, the Kron-based reduced network is achieved by reducing the set of nodes $\mathcal{R}=\{i\in\mathcal{V}|I^{\rm agg}_{i}=\mathbf{0}\}$. Consider assignment of nodes with nonzero injections results in $V^{\rm agg}\neq V$. This, in turn, results in an error in voltages between nodes in the reduced and full networks. Consequently, in the next section, we aim to find the optimal assignment matrix to maximize number of reduced nodes and accuracy simultaneously.
@@ III Optimal Kron-based Reduction of Networks
The choice of $A$ determines how injections are aggregated and how voltages are represented, and thus directly affects both the reduction level and the accuracy of the reduced model. In general, greater reduction introduces larger errors. The objective is to find the assignment matrix $A$ by finding a desired balance between reduction and accuracy. One way to achieve this is to cast the search for $A$ as an optimization problem, where the assignment rules – phase availability, connectivity, and uniqueness – are enforced as constraints, as was done in Opti-KRON [9] for single phase networks.
In this section, we extend the formulation to three-phase feeders. The core optimization framework follows the single-phase structure but is reformulated to handle three-phase feeders. Let $\mathcal{L}$ denote a library of operating scenarios given from AC load flow solutions on the full network. For each scenario $l\in\mathcal{L}$, $\hat{V}_{l}\in\mathbb{C}^{3n}$ and $\hat{I}_{l}\in\mathbb{C}^{3n}$ are the three-phase nodal voltage and current injection vectors of the full network that satisfy $Y\hat{V}_{l}=\hat{I}_{l}$. The voltage of the network under aggregated current is denoted by $V_{l}$, such that $YV_{l}=A\hat{I}_{l}$. In this paper, error is defined as the difference between the nodal voltage in the full and in the reduced network. In the reduced network, super-node voltages are obtained through (12), while each reduced node is assigned the voltage of its associated super-node. Accordingly, for $l\in\mathcal{L}$, error is computed as
${E_{l}}=\overbrace{(A\otimes\mathbf{I}_{3}){\rm diag}\{\hat{V}_{l}\}}^{\text{Full}}-\overbrace{{\rm diag}\{V_{l}\}(A\otimes\mathbf{I}_{3})}^{\text{ Reduced}}\in\mathbb{C}^{n\times n}.$ (13)
@@
$\text{MICE}=\left\|\vec{e}_{i}^{T}E_{l}\right\|_{\infty}.$ (14)
Note that $\vec{e}_{i}$ represent the $i^{\rm th}$ standard basis vector. Using MICE as the measure of accuracy, we formulate the three-phase Opti-KRON problem, to find the assignment matrix that as follows.
@@
The objective function(15a) minimizes the sum MICE across all scenarios, while simultaneously minimizing the number of super-nodes. The trade-off is tuned by the parameter $\alpha$. The KCL equation is enforced in (15c). Voltage magnitude deviations are constrained in (15d) to lie within a tolerance $\bar{E}$. Unique assignments are enforced by (15e) to ensure each reduced node maps to exactly one super-node. Equation (15f) allows assignments only to super-nodes. To keep the clusters connected, (15g) restricts clustering to adjacent nodes through the adjacency matrix $\Lambda$, while (15h) enforces availability of phases of reduced nodes at super-nodes. The constraint (15i) bounds the number of reduced nodes, to speed-up the solve times. Finally, (15j) declares all assignment variables as binary.
To solve (15), we follow the methodology adapted from [9], which addressed the same challenges:
@@
    The problem is a mixed-integer nonlinear program (MINLP), due to the nonlinear voltage constraints (15b) and (15d). This can be relaxed to a mixed-integer linear program (MILP) by introducing auxiliary variables and Big M method.
@@
To iteratively solve (15), the adjacency matrix must be updated after each iteration to represent new admissible assignments. After each iteration $t$, based on how reduced nodes are assigned, we update the adjacency matrix to model connections among super-nodes as
@@
where, $A^{*}$ represents the optimal assignment matrix obtained at iteration $t$, and $\odot$ denotes element-wise multiplication. It connects each super-node to the neighbors of itself and its assigned nodes. If the original network is radial, $\Lambda$ remains radial throughout. The number of reduced nodes $n_{r}^{(t)}$ and super-nodes $n_{s}^{(q,t)}$ at iteration $t$ are both functions of the reduction limit $q$. The number of nodes that can be reduced in iteration $t$ is bounded by $n_{r}^{(t)}\leq q$, and the number of remaining super-nodes evolves as
@@
Since $\Lambda^{t}$ is radial, the number of lines at each iteration is $n_{s}^{(q,t)}-1$. Accordingly, the number of distinct feasible reduction scenarios that is restricted by (15g) equals $2(n_{s}^{(q,t)}-1)$, i.e., two scenarios per available line (for every line, load could move across in either direction). Here, by distinct feasible reduction scenarios, we mean the total number of pairs $(i,j)$ that satisfy (15g).
@@ IV Exhaustive Search for Reduction
Building on the MILP formulation of Section III, which optimally determines reduction assignments through an MILP, we now introduce a complementary approach designed for parallel execution on GPUs. Motivated by Lemma 1, reducing one node per iteration let us evaluate each distinct reduction scenario independently and in parallel. Additionally, while solving the MILP requires decomposing complex variables into their real and imaginary parts, the exhaustive search approach operates directly in the complex domain and enables exact computation of voltage deviations. Furthermore, unlike the MILP, which minimizes a linear objective, the exhaustive-search framework allows direct evaluation of nonlinear error measures. In particular, objectives based on voltage magnitudes, while inherently nonconvex, can be assessed exactly for each candidate reduction. Accordingly, we adjust the objective of each iteration to minimize the sum of MICE (SMICE), where error is measured as voltage magnitude deviation. At each iteration, we are find the assignment such that
@@
We initialize the assignment set such that every node is a super-node and is assigned to itself. At each iteration, the set of candidate assignments must satisfy phase availability and connectivity constraints. The algorithm then evaluates the error of every potential assignment, selects the candidate with the minimum sum of MICE (SMICE) across all loading scenarios, and proceeds with updated adjacency, current vector, and assignment set. For all iterations, $Y$ remains fixed, while the current vector $I^{c}$ is updated for each assignment. Solving $I^{c}=YV^{c}$ results in the same super-nodes voltage that Kron reduction would yield. After each solve, reduced nodes are assigned the voltage of their super-node, both for new and previously reduced nodes. The parallelizable part of the algorithm is the evaluation of candidate assignments, where each candidate assignment is evaluated independently with no data dependencies. This independency makes this block suitable for parallel computation of the resulting error of each candidate assignments ($\rm SMICE$}), which, in turn, makes the proposed approach scalable. However, the Kron-based reduction of a radial network generates smaller, but dense meshed network [2, 9]. To recover a radial three-phase distribution network, we now review and extend the radialization approach introduced in [9].
@@ V Three-phase Radialization
Kron-based reduced networks are prone to being meshed and densely connected. This structure makes us unable to use power flow algorithms optimized for three-phase radial networks. Density of the reduced networks increases the computational burden of subsequent optimization problems. In [9], radialization was presented to find a radial equivalent of a single-phase Kron reduced network. Here, we extend that approach to three-phase networks.
The topology of a Kron reduced graph can be realized through its admittance matrix from (9). In single-phase networks, any nonzero entry ${Y_{\rm Kron}}_{i,j}$ indicates a direct connection between nodes $i$ and $j$. For three-phase feeders, we extend this definition: a connection between nodes $i$ and $j$ exists if the corresponding $3\times 3$ block in $Y_{\rm Kron}$} contains at least one nonzero entry. In the following, we first demonstrate that topology of a three-phase Kron reduced network can be found through its Kron reduced graph Laplacian as Lemma 2. Then, we leverage this observation to extend single-phase radialization to three-phase cases.
@@
From (23), we can infer that the only part of the reduced network that may have different topology from the full network is the connection among nodes in $\mathcal{K}_{a}$. Since the original admittance matrix $Y$ corresponds to a radial network and no two distinct neighbors of $r$ are directly connected, $Y_{\mathcal{K}_{a},\mathcal{K}_{a}}$ is block diagonal of size $3|\mathcal{K}_{a}|\times 3|\mathcal{K}_{a}|$. On the other hand, every node in $\mathcal{K}_{a}$ is adjacent to $r$, which implies that all the elements in $Y_{\mathcal{K}_{a},r}\in\mathbb{C}^{3|\mathcal{K}_{a}|\times 3}$, $Y_{r,\mathcal{K}_{a}}\in\mathbb{C}^{3\times 3|\mathcal{K}_{a}|}$, and $Y_{r,r}^{+}\in\mathbb{C}^{3\times 3}$ have non-zeros values. Therefore, $Y_{\mathcal{K}_{a},r}\,Y_{r,r}^{+}\,Y_{r,\mathcal{K}_{a}}$ and, in turn, the lower right block of (23) yields a fully dense sub-matrix with all the values being non-zeros. This implies that, after eliminating node $r$, all nodes in $\mathcal{K}_{a}$ become mutually connected, while the connections among nodes in $\mathcal{K}_{f}$, remain unchanged. The same structural effect occurs in the single-phase Kron reduction [9]. Hence, Kron reduction of a three-phase network and a single-phase network with the same topology, e.g., its graph Laplacian, results in reduced networks with identical topologies. ∎
Fig. 4 demonstrates different stage of Kron reduction.
@@
Reference [11] proves that Kron reduction of radial networks is an edge-disjoint of maximal cliques. Lemma 2 extends this property to three-phase networks. A clique is a sub-graph that all the nodes in the sub-graph are mutually connected. When a clique is not included in any larger clique it is call a maximal clique. Any maximal cliques with three or more nodes creates meshes in Kron reduced networks, which is the consequence of eliminating nodes with degree $\geq 3$. These properties are identical between single and three-phase networks. Therefore, we summarize the radialization procedure of [9], which can be extended directly into three-phase networks.

### v2

@@ I Introduction
The rapid integration of distributed energy resources (DERs) and advanced control schemes has significantly increased the complexity of distribution networks, posing challenges for optimization and real-time decision-making [undef]. Detailed network models with thousands of nodes are computationally intensive and are impractical for large-scale optimal power flow and control problems. Consequently, research on network reduction has emerged, where proposed network reduction tools seek to simplify network models while preserving essential electrical characteristics.
@@
One network reduction technique is bus elimination via Kron reduction [undefa], where selected nodes are eliminated via the Schur complement of the nodal admittance matrix, resulting in an equivalent network that preserves the impedance seen between kept nodes. However, the set of reducible nodes is restricted to those with zero current injection. References [undefb, undefc] develop network reduction methods for single-phase distribution systems. In [undefb], the network parameters of a reduced model are updated through optimization, whereas [undefc] reduces the feeder by recursively merging parts of the radial graph into equivalent nodes. However, both approaches are formulated for single-phase systems and therefore do not extend to real-world unbalanced three-phase networks.
For unbalanced three-phase distribution feeders, several reduction approaches have been proposed. In [undefd, undefe], reduced feeders are constructed by eliminating non-critical buses from the feeder impedance matrices and reconstructing the corresponding network elements for the kept nodes. Similarly, [undeff, undefg, undefh] construct reduced networks by replacing selected parts of the feeder with simplified line-based representations. In [undeff, undefh], the reduced network is represented by a single segment between two remaining nodes, whereas [undefg] replaces multiple sectors, each with one line. A graph-based reduction approach is presented in [undefi], where a minimum-spanning-tree search is used to identify feeder sections for aggregation into a smaller reduced network. Application-specific reduced feeders are developed in [undefj, undefk, undefl], where selected buses are retained and the remaining network is reduced for power flow analysis, real-time simulation, or risk-aware islanding studies. The existing network reduction methods discussed so far assume a predefined set of nodes to be retained in the reduced network. However, the quality of the reduced network depends strongly on that choice. In general, more aggressive network reduction leads to larger approximation errors. Therefore, incorporating an error metric into the reduction process to determine the set of kept nodes and balance reduction level against accuracy is important. However, this aspect is lacking in existing network reduction methods for three-phase feeders.
@@
In prior works [undefm] and [undefn], an optimal Kron-based reduction of networks (Opti-KRON) framework has been developed for single-phase balanced feeders, where the reduced networks do not rely on any predefined set of nodes and explicitly trade off reduction level against accuracy via tunable bounds on voltage magnitude error. A mixed-integer linear program (MILP) was proposed to identify the optimal set of nodes and clusters that maximize reduction while minimizing voltage deviations across representative loading scenarios. In addition, because Kron reduction can yield dense, meshed networks [undefa][undefm], the framework introduced a radialization step to ensure the reduced network remains radial. However, the previous Opti-KRON formulation is limited to single-phase networks. To the best of our knowledge, the literature lacks a scalable approach that computes optimal network reductions for unbalanced three-phase distribution feeders.
@@
In this paper, we extend the Opti-KRON framework of [undefm] to unbalanced three-phase distribution networks. We develop a three-phase Kron-based reduction and radialization procedure. To formulate the optimal reduction, we first generalize the single-phase Opti-KRON model to three-phase feeders. Furthermore, we introduce an exhaustive-search algorithm, with the MILP formulation serving as a methodological baseline. The proposed exhaustive search offers several advantages. First, it operates directly on complex values without requiring decomposition into rectangular coordinates. Second, it can evaluate nonconvex objective functions. Third, its structure supports parallel computation on GPUs. Consequently, the proposed method achieves significant computational speedups on large-scale feeders while increasing reduction accuracy. This paper makes the following contributions:
@@
The remainder of this paper is organized as follows. In Section II, the three-phase network model and Kron reduction are summarized. The concept of Kron-based network reduction is introduced in Section III. Section IV introduces our exhaustive-search framework for optimal Kron-based reduction. We then describe the radialization step in Section V. Experimental results are presented in Section VI. We conclude in Section VII with a summary and future directions.
@@ II-A Network Representation
Consider a distribution network as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n=\lvert\mathcal{V}\rvert$ nodes and $m=\lvert\mathcal{E}\rvert$ branches. We define the network topology via the adjacency matrix $\Lambda\in\{0,1\}^{n\times n}$, where $\Lambda_{ij}=1$, if a branch connects nodes $i$ and $j$, and $\Lambda_{ij}=0$ otherwise. Each node $i$ carries a subset of phases $\phi_{i}\subseteq\{a,b,c\}$. Each node $i$ is associated with a three-phase voltage vector $V_{i}\in\mathbb{C}^{3}$ and current injection vector $I_{i}\in\mathbb{C}^{3}$. Complex nodal voltage and current injection vectors are defined as
@@ II-B Kron Reduction of Three-Phase Networks
The Kron reduction of $Y$ is the Schur complement of $Y_{\mathcal{R}\mathcal{R}}$ [undefa]. Thus, the Kron reduced admittance matrix, denoted as $Y_{\rm Kron}$, is given by
@@
where $\{\cdot\}^{+}$ represents the Moore-Penrose Pseudoinverse  [^2] The pseudo-inverse is utilized to account for zero rows and columns associated with missing phases during elimination. and
@@ II-C Clustering for Kron-based Reduction
As shown in (8) and (10), Kron reduction preserves voltage profile when the reduced nodes have zero current injection. To enable reduction beyond strictly zero-injection nodes, we introduce an assignment scheme, where each reduced node moves its current injection to a kept adjacent node. We refer to the kept nodes $\mathcal{K}$ as super-nodes; each super-node contains its own current injection together with those of the reduced nodes assigned to it. Each super-node $k\in\mathcal{K}$ represents a cluster consisting of itself and its assigned children-nodes. These assignments, however, must obey several rules:
@@
    Phase availability: a reduced node $i$ can only be assigned to a super-node $k$ if $\phi_{i}\subseteq\phi_{k}$.
@@
Now, the Kron-based reduced network is achieved by reducing the set of nodes $\mathcal{R}=\{i\in\mathcal{V}|I^{\rm agg}_{i}=\mathbf{0}\}$. The assignment of nodes with nonzero injections generally leads to $V^{\rm agg}\neq V$. This, in turn, results in an error in voltages between nodes in the reduced and full networks. Consequently, in the next section, we aim to find the optimal assignment matrix to maximize number of reduced nodes and accuracy simultaneously.
@@ III Optimal Kron-based Reduction of Networks
The choice of $A$ determines how injections are aggregated and how voltages are represented, and thus directly affects both the reduction level and the accuracy of the reduced model. In general, greater reduction introduces larger errors. The objective is to find the assignment matrix $A$ by finding a desired balance between reduction and accuracy. One way to achieve this is to cast the search for $A$ as an optimization problem, where the assignment rules – phase availability, connectivity, and uniqueness – are enforced as constraints, as was done in Opti-KRON [undefm] for single phase networks.
In this section, we extend the formulation to three-phase feeders. The core optimization framework follows the single-phase structure but is reformulated to handle three-phase feeders. Let $\mathcal{L}$ denote a library of operating scenarios given from AC load flow solutions on the full network. For each scenario $l\in\mathcal{L}$, $\hat{V}_{l}\in\mathbb{C}^{3n}$ and $\hat{I}_{l}\in\mathbb{C}^{3n}$ are the three-phase nodal voltage and current injection vectors of the full network that satisfy $Y\hat{V}_{l}=\hat{I}_{l}$. The voltage of the network under aggregated current is denoted by $V_{l}$, such that $YV_{l}=A\hat{I}_{l}$. In this paper, error is defined as the difference between the nodal voltage in the full and reduced networks. In the reduced network, super-node voltages are obtained through (12), while each reduced node is assigned the voltage of its associated super-node. Accordingly, for $l\in\mathcal{L}$, error is computed as
${E_{l}}=\overbrace{(A\otimes\mathbf{I}_{3}){\rm diag}\{\hat{V}_{l}\}}^{\text{Full}}-\overbrace{{\rm diag}\{V_{l}\}(A\otimes\mathbf{I}_{3})}^{\text{ Reduced}}\in\mathbb{C}^{3n\times 3n}.$ (13)
@@
$\text{MICE}_{i,l}=\left\|\vec{e}_{i}^{T}E_{l}\right\|_{\infty}.$ (14)
Note that $\vec{e}_{i}$ represent the $i^{\rm th}$ standard basis vector. Using MICE as the measure of accuracy, we formulate the three-phase Opti-KRON problem to find the assignment matrix as follows.
@@
The objective function (15a) minimizes the sum of MICE values across all scenarios, while simultaneously minimizing the number of super-nodes. The trade-off is tuned by the parameter $\alpha$. The KCL equation is enforced in (15c). Voltage magnitude deviations are constrained in (15d) to lie within a tolerance $\bar{E}$. Unique assignments are enforced by (15e) to ensure each reduced node maps to exactly one super-node. Equation (15f) allows assignments only to super-nodes. To keep the clusters connected, (15g) restricts clustering to adjacent nodes through the adjacency matrix $\Lambda$, while (15h) enforces availability of the reduced node phases at the super-nodes. The constraint (15i) bounds the number of reduced nodes, to speed-up the solve times. Finally, (15j) declares all assignment variables as binary.
To solve (15), we follow the methodology adapted from [undefm], which addressed the same challenges:
@@
    The problem is a mixed-integer nonlinear program (MINLP), due to the nonlinear voltage constraints (15b) and (15d). This can be relaxed to a mixed-integer linear program (MILP) by introducing auxiliary variables and Big M method. Accordingly, the resulting MILP should be interpreted as a relaxation of the original nonconvex MINLP: feasibility is imposed through linearized voltage magnitude error constraints, while the exact voltage errors are evaluated a posteriori.
@@
To iteratively solve (15), the adjacency matrix must be updated after each iteration to reflect newly admissible assignments. After each iteration $t$, based on how reduced nodes are assigned, we update the adjacency matrix to model connections among super-nodes as
@@
where, $A^{*}$ represents the optimal assignment matrix obtained at iteration $t$, and $\odot$ denotes element-wise multiplication.  (16) connects each super-node to the neighbors of itself and its assigned nodes. If the original network is radial, $\Lambda$ remains radial throughout. The number of reduced nodes $n_{r}^{(t)}$ and super-nodes $n_{s}^{(q,t)}$ at iteration $t$ are both functions of the reduction limit $q$. The number of nodes that can be reduced in iteration $t$ is bounded by $n_{r}^{(t)}\leq q$, and the number of remaining super-nodes evolves as
@@
Since $\Lambda^{t}$ is radial, the number of lines at each iteration is $n_{s}^{(q,t)}-1$. Accordingly, the number of distinct feasible reduction scenarios restricted by (15g) equals $2(n_{s}^{(q,t)}-1)$, i.e., two scenarios per available line (for every line, the load could move across in either direction). Here, by distinct feasible reduction scenarios, we mean the total number of pairs $(i,j)$ that satisfy (15g).
@@ IV Exhaustive Search for Reduction
Building on the MILP formulation of Section III, which optimally determines reduction assignments, we now introduce a complementary approach designed for parallel execution on GPUs. Motivated by Lemma 1, reducing one node per iteration allows us to evaluate each distinct reduction scenario independently and in parallel. Additionally, while solving the MILP requires decomposing complex variables into their real and imaginary parts, the exhaustive search approach operates directly in the complex domain and enables exact computation of voltage deviations. Furthermore, unlike the MILP, which minimizes a linear objective, the exhaustive-search framework allows direct evaluation of nonlinear error measures. In particular, objectives based on voltage magnitudes, while inherently nonconvex, can be assessed exactly for each candidate reduction. Accordingly, we adjust the objective of each iteration to minimize the sum of MICE (SMICE), where error is measured as voltage magnitude deviation. At each iteration, we find the assignment such that
@@
We initialize the assignment set such that every node is a super-node and is assigned to itself. At each iteration, the set of candidate assignments must satisfy phase availability and connectivity constraints. The algorithm then evaluates the error of every potential assignment, selects the candidate with the minimum sum of MICE (SMICE) across all loading scenarios, and proceeds with updated adjacency, current vector, and assignment set. For all iterations, $Y$ remains fixed, while the current vector $I^{c}$ is updated for each assignment. Solving $I^{c}=YV^{c}$ results in the same super-node voltages that Kron reduction would yield. After each solve, reduced nodes are assigned the voltage of their super-node, both for new and previously reduced nodes. The parallelizable part of the algorithm is the evaluation of candidate assignments, where each candidate assignment is evaluated independently with no data dependencies. This independence makes this block suitable for parallel computation of the resulting error of each candidate assignment ($\rm SMICE$}), which, in turn, makes the proposed approach scalable. However, the Kron-based reduction of a radial network generates a smaller but dense meshed network [undefa, undefm]. To recover a radial three-phase distribution network, we now review and extend the radialization approach introduced in [undefm].
@@ V Three-phase Radialization
Kron-based reduced networks are prone to being meshed and densely connected. This structure prevents the use of power flow algorithms optimized for three-phase radial networks. The density of the reduced networks increases the computational burden of subsequent optimization problems. In [undefm], radialization was presented to find a radial equivalent of a single-phase Kron reduced network. Here, we extend that approach to three-phase networks.
The topology of a Kron reduced graph can be realized through its admittance matrix from (9). In single-phase networks, any nonzero entry ${Y_{\rm Kron}}_{i,j}$ indicates a direct connection between nodes $i$ and $j$. For three-phase feeders, we extend this definition: a connection between nodes $i$ and $j$ exists if the corresponding $3\times 3$ block in $Y_{\rm Kron}$} contains at least one nonzero entry. In the following, we first demonstrate that the topology of a three-phase Kron reduced network can be found through its Kron reduced graph Laplacian as Lemma 2. Then, we leverage this observation to extend single-phase radialization to three-phase cases.
@@
From (23), we can infer that the only part of the reduced network that may have different topology from the full network is the connection among nodes in $\mathcal{K}_{a}$. Since the original admittance matrix $Y$ corresponds to a radial network and no two distinct neighbors of $r$ are directly connected, $Y_{\mathcal{K}_{a},\mathcal{K}_{a}}$ is block diagonal of size $3|\mathcal{K}_{a}|\times 3|\mathcal{K}_{a}|$. On the other hand, every node in $\mathcal{K}_{a}$ is adjacent to $r$, which implies that all the elements in $Y_{\mathcal{K}_{a},r}\in\mathbb{C}^{3|\mathcal{K}_{a}|\times 3}$, $Y_{r,\mathcal{K}_{a}}\in\mathbb{C}^{3\times 3|\mathcal{K}_{a}|}$, and $Y_{r,r}^{+}\in\mathbb{C}^{3\times 3}$ have non-zero values. Therefore, $Y_{\mathcal{K}_{a},r}\,Y_{r,r}^{+}\,Y_{r,\mathcal{K}_{a}}$ and, in turn, the lower right block of (23) yields a fully dense sub-matrix with all the values being non-zeros. This implies that, after eliminating node $r$, all nodes in $\mathcal{K}_{a}$ become mutually connected, while the connections among nodes in $\mathcal{K}_{f}$, remain unchanged. The same structural effect occurs in the single-phase Kron reduction [undefm]. Hence, Kron reduction of a three-phase network and a single-phase network with the same topology, e.g., its graph Laplacian, results in reduced networks with identical topologies. ∎
Fig. 4 demonstrates different stages of Kron reduction.
@@
Reference [undefo] proves that Kron reduction of radial networks is an edge-disjoint union of maximal cliques. Lemma 2 extends this property to three-phase networks. A clique is a sub-graph that all the nodes in the sub-graph are mutually connected. When a clique is not included in any larger clique, it is called a maximal clique. Any maximal cliques with three or more nodes creates meshes in Kron reduced networks, which is the consequence of eliminating nodes with degree $\geq 3$. These properties are identical between single and three-phase networks. Therefore, we summarize the radialization procedure of [undefm], which can be extended directly into three-phase networks.